# Inexact subgradient algorithm with a non-asymptotic convergence guarantee for copositive programming problems

### v1

@@ 3 An inexact subgradient algorithm for convex semi-infinite programming problems
    For every $\bm{\delta}\in\Delta$, $g(\cdot;\bm{\delta})\colon V\to(-\infty,+\infty]$ is a closed convex function that is subdifferentiable on $S$. Moreover, there exists a positive constant $L_{g}$ such that $\lVert d\rVert\leq L_{g}$ for all $x\in S$ and $d\in\partial_{x}g(x;\bm{\delta})$.
@@
where (a) follows from the update rule in (i) of Algorithm 1, (b) holds by $x^{*}\in S$ and the nonexpansiveness of the projection $P_{S}$, (c) follows from $f^{*}=f(x^{*})$ and $d_{k}\in\partial f(x_{k})$, (d) is a consequence of the contradictory assumption, and (e) follows from $\lVert d_{k}\rVert\leq L$. In addition, for every $k\in[N]\setminus I_{N}$, we have
@@
holds at each iteration. If we know a way to evaluate the exact value of $G(x)$, we can set $\alpha=0$. Setting $\alpha$ to a positive value allows us to compute $G(x)$ approximately. By substituting $\alpha\epsilon$ for $\xi_{k}$ in (3.4), Lemma 3.3 directly implies the following theorem.
@@
Table 1: This table summarizes the number of iterations required to obtain $\bar{x}\in S$ such that $f(\bar{x})\leq f^{*}+\epsilon$ and $G(\bar{x})\leq\epsilon$ when we solve Problem (3.1) using each algorithm. Here, $S$ is assumed to be a compact set of diameter $D$. Bregman distances are utilized in the algorithms of [5] and [55], but we show the number of iterations when the Bregman distance is Euclidean. For the algorithm in [55], the inequality $G(\bar{x})\leq\epsilon$ holds in expectation. In addition, the algorithm in [55] yields $\bar{x}\in S$ such that $f(\bar{x})\leq f^{*}+\frac{3}{7}\epsilon$, while Algorithm 1 yields $\bar{x}\in S$ such that $f(\bar{x})\leq f^{*}+\frac{1}{1+\alpha}\epsilon$. The number of iterations in the algorithm of [5] is obtained by taking the limit as $\beta\to 1$, where $\beta<1$ is a parameter appearing in [5, Corollary 2.1].
@@ 4.1 General copositive programming problems
with the coefficient matrix $\bm{Q}\in\mathcal{S}^{n}$ given by $\bm{A}_{0}+\sum_{i=1}^{m}x_{i}\bm{A}_{i}$. In the following subsections, we give an overview of several approaches to solving a standard quadratic programming problem of the form (4.7), both exactly (the case where the violation parameter $\alpha$ equals 0) and inexactly (the case where $\alpha>0$). An exact method is presented in Section 4.1.1. Inexact methods are presented in Sections 4.1.2, 4.1.3, and 4.1.4. Whereas the exact method requires external solvers to solve a mixed-integer linear programming problem, only a finite number of function evaluations are performed in the inexact methods. In Sections 4.1.2, 4.1.3, and 4.1.4, we only consider the case where $\alpha=1$. This is without loss of generality because for general $\alpha>0$, we regard $\alpha\epsilon$ as a new $\epsilon$.
In the following methods, we frequently use a lower bound for $\gamma(\bm{Q})$. The choice of the lower bound is arbitrary, and there are many alternatives that are summarized in the literature [11]. In our study, because we need to calculate the lower bound per iteration, it might be preferable to adopt a computationally inexpensive one. As such a bound, under the convention that $1/0=\infty$, $a+\infty=\infty$ for any $a\in\mathbb{R}\cup\{\infty\}$, and $1/\infty=0$, Bomze, Locatelli, and Tardella [11] provided the following lower bound
$\underline{\gamma}(\bm{Q})\coloneqq\min_{1\leq i\leq j\leq n}Q_{ij}+\frac{1}{\displaystyle\sum_{k=1}^{n}\frac{1}{Q_{kk}-{\displaystyle\min_{1\leq i\leq j\leq n}Q_{ij}}}}$
for $\gamma(\bm{Q})$ and showed its superiority over other simple lower bounds.
@@ 4.1.1 Exact method through mixed-integer linear programming
To solve the standard quadratic programming problem in (4.7) exactly, we can utilize the mixed-integer linear programming reformulations proposed by Gondzio and Yıldırım [22]. In their paper, various mixed-integer linear programming reformulations of standard quadratic programming problems are presented. Among them, we adopt the following reformulation based on [22, Proposition 2]:*1*1*1Using their terminology, we adopt formulation (MILP2) with the lower bound $\ell_{1}(\bm{Q})$. They claim that this reformulation is suitable for large-scale instances because of its robust practical performance [22, page 319].
@@ 4.1.2 Inexact deterministic method through a regular grid of the standard simplex
for a positive integer $r$. From [8, Theorem 3.2] (see also [41, Theorem 2]), we have
@@
which is polynomial in $n$ if we regard $r$ as a constant.
@@ 4.1.3 Inexact randomized method through uniform sampling from the standard simplex
The second inexact method for solving the standard quadratic programming problem is to discretize the standard simplex $\Delta^{n-1}$ by uniformly sampling from it. Let $\bm{\delta}_{1},\dots,\bm{\delta}_{M}$ be a collection of independent and identically distributed random vectors uniformly distributed on $\Delta^{n-1}$. In Proposition 4.2 shown below, we provide a lower bound for the number $M$ of samples sufficient to achieve
$\min_{1\leq i\leq M}\bm{\delta}_{i}^{\top}\bm{Q}\bm{\delta}_{i}-\gamma(\bm{Q})\leq\epsilon.$ (4.14)
For notational convenience, we define
$m(\rho,\phi)\coloneqq\frac{\log\phi}{\log(1-\rho)}$
for $\rho,\phi\in(0,1)$. In addition, throughout this subsubsection, $P$ denotes the uniform distribution on $\Delta^{n-1}$.
@@
By [29, Lemma 6.21], the function $\operatorname*{vol}(B(\bm{x},r)\cap\Delta^{n-1})^{n-1}$ with respect to $\bm{x}\in\Delta^{n-1}$ is concave, so it achieves the minimum at an extreme point of $\Delta^{n-1}$. By the symmetry of $\Delta^{n-1}$, it does so at $\bm{x}=\bm{e}_{1}$.
@@
Let $\bm{\delta}_{1},\dots,\bm{\delta}_{M}$ be a collection of independent and identically distributed random vectors uniformly distributed on $\Delta^{n-1}$. For $\epsilon\in(0,\sqrt{2}K(\bm{Q})]$ and $\phi\in(0,1)$, if the number $M$ of samples satisfies
@@ 4.2 Testing complete positivity of a matrix
where we use [27, Equation (6.16)] to derive the first inequality and we use (4.32) to derive the second inequality. Using this inequality, the Cauchy–Schwarz inequality, and $\lVert\bm{C}\rVert_{\rm F}=1$, we have
@@
iterations are necessary to test non-complete positivity by using Theorem 4.8.
@@ 4.3 The case of symmetric cones beyond nonnegative orthants
In what follows, we explore a potential approach to deriving a global solution to Problem (4.34). It follows from [18, Section 3] that the optimal value of (4.34) is equal to that of the following copositive programming problem with a scalar variable:
@@ 5.1 Comparison among the methods for solving standard quadratic programming problems
Table 2: The results of solving standard quadratic programming problems by using MILP, Grid, and SUnif. The average objective values obtained by solving standard quadratic programming problems with each method are listed in the “Obj. val.” column and the average computational time required to solve them is listed in the “Time [s]” column. The “Dev. from MILP” column denotes the difference between each average objective value and that obtained by using MILP for the same $n$. The “MM” column denotes the average number $M$ of points $\bm{\delta}_{1},\dots,\bm{\delta}_{M}\in\Delta^{n-1}$ that are used to calculate an approximate optimal value $\min_{1\leq i\leq M}\bm{\delta}_{i}^{\top}\bm{Q}\bm{\delta}_{i}$ for a standard quadratic programming problem of the form (4.7).
@@
0.1 >>3600 >>3600 >>3600 >>3600
SUnif 1, 0.1 >>3600 >>3600 >>3600 >>3600
@@
0.1 >>3600 >>3600 >>3600 >>3600
SUnif 1, 0.1 >>3600 >>3600 >>3600 >>3600
@@
0.1 >>3600 >>3600 >>3600 >>3600
SUnif 1, 0.1 >>3600 >>3600 >>3600 >>3600
@@
0.1 >>3600 >>3600 >>3600
SUnif 1, 0.1 >>3600 >>3600 >>3600
@@
Second, to see the effectiveness of the randomized method introduced in Section 4.1.4 (denoted GUnif), we compare it with the other randomized method, SUnif, in terms of solution accuracy and computational time. Recall that in SUnif we sample from the standard simplex $\Delta^{n-1}$, whereas in GUnif we sample points from the regular grid $\Delta_{r}^{n-1}$ for some positive integer $r$. For each $n$, we solved only one of the ten instances. We set $\epsilon$ to $1$ or $0.1$, and for each $(n,\epsilon)$ we set $r$ as in the previous experiment and choose the number $M$ of samples to be the smallest integer greater than or equal to $c\lvert\Delta_{r}^{n-1}\rvert$, where $c\in\{$5\text{\times}{10}^{-1}$,$1\text{\times}{10}^{-1}$,$1\text{\times}{10}^{-2}$,$1\text{\times}{10}^{-3}$,$1\text{\times}{10}^{-4}$\$\}. We sampled $M$ points $\bm{\delta}_{1},\dots,\bm{\delta}_{M}$ from the standard simplex $\Delta^{n-1}$ when we used SUnif, whereas we sampled from the grid $\Delta_{r}^{n-1}$ when we used GUnif. Then we calculated the value $\min_{1\leq i\leq M}\bm{\delta}_{i}^{\top}\bm{Q}\bm{\delta}_{i}$.
[…]
We observed that, except for the two cases where the numbers of samples were less than $10$, the objective values obtained by using GUnif were always better than those obtained with SUnif. However, in many cases, especially in Table 4, the computational time for GUnif was longer than that for SUnif. The MATLAB Profiler showed that our implementation of GUnif spent most time in the built-in function sort, which we used to implement uniform sampling from the regular grid. A more efficient implementation of GUnif would shorten its computational time.
5.2 Solving general copositive programming problems
@@
$1\text{\times}{10}^{-3}$ 22 $2.61\text{\times}{10}^{-01}$ $3.84\text{\times}{10}^{-02}$ $7.13\text{\times}{10}^{-01}$ $2.1\text{\times}{10}^{-03}$ $1.6\text{\times}{10}^{-03}$
$\text{\times}{10}^{-4} 11$ $1.75\text{\times}{10}^{-01}$ $1.33\text{\times}{10}^{-02}$ $7.98\text{\times}{10}^{-01}$ $1.0\text{\times}{10}^{-03}$ $1.5\text{\times}{10}^{-03}$
@@
$1\text{\times}{10}^{-3}$ 66 $3.00\text{\times}{10}^{-03}$ $1.46\text{\times}{10}^{-02}$ $9.39\text{\times}{10}^{-01}$ $1.4\text{\times}{10}^{-03}$ $1.7\text{\times}{10}^{-03}$
$\text{\times}{10}^{-4} 11$ $1.99\text{\times}{10}^{-01}$ $3.02\text{\times}{10}^{-02}$ $7.43\text{\times}{10}^{-01}$ $4.9\text{\times}{10}^{-03}$ $2.2\text{\times}{10}^{-03}$
@@
$1\text{\times}{10}^{-3}$ $126$ $8.04\text{\times}{10}^{-01}$ $8.11\text{\times}{10}^{-03}$ $1.94\text{\times}{10}^{-01}$ $8.7\text{\times}{10}^{-03}$ $3.2\text{\times}{10}^{-03}$
$1\text{\times}{10}^{-4}$ $13$ $6.66\text{\times}{10}^{-01}$ $7.31\text{\times}{10}^{-03}$ $3.31\text{\times}{10}^{-01}$ $2.0\text{\times}{10}^{-03}$ $1.7\text{\times}{10}^{-03}$
@@
$1\text{\times}{10}^{-3}$ $501$ $8.04\text{\times}{10}^{-01}$ $4.49\text{\times}{10}^{-03}$ $1.89\text{\times}{10}^{-01}$ $3.6\text{\times}{10}^{-02}$ $1.5\text{\times}{10}^{-02}$
$1\text{\times}{10}^{-4}$ $51$ $7.53\text{\times}{10}^{-01}$ $4.19\text{\times}{10}^{-03}$ $2.40\text{\times}{10}^{-01}$ $6.0\text{\times}{10}^{-03}$ $4.2\text{\times}{10}^{-03}$

### v2

@@ 3 An inexact subgradient algorithm for convex semi-infinite programming problems
    For every $\bm{\delta}\in\Delta$, $g(\cdot;\bm{\delta})\colon V\to(-\infty,+\infty]$ is a closed convex function that is subdifferentiable on $S$. Moreover, there exists a positive constant $L_{g}$ such that $\lVert d\rVert\leq L_{g}$ for all $x\in S$, $\bm{\delta}\in\Delta$, and $d\in\partial_{x}g(x;\bm{\delta})$.
@@
where (a) follows from the update rule in (i) of Algorithm 1, (b) holds by $x^{*}\in S$ and the nonexpansiveness of the projection $P_{S}$, (c) follows from $f^{*}=f(x^{*})$ and $d_{k}\in\partial f(x_{k})$, (d) is a consequence of the assumption for contradiction, and (e) follows from $\lVert d_{k}\rVert\leq L$. In addition, for every $k\in[N]\setminus I_{N}$, we have
@@
holds at each iteration. If we know a way to evaluate the exact value of $G(x)$, we can set $\alpha=0$. Setting $\alpha$ to a positive value allows us to compute $G(x)$ approximately. Since $\xi_{k^{*}}$ is also bounded by $\alpha\epsilon$, Lemma 3.3 directly implies the following theorem.
@@
Table 1: This table summarizes the number of iterations required to obtain $\bar{x}\in S$ such that $f(\bar{x})\leq f^{*}+\epsilon$ and $G(\bar{x})\leq\epsilon$ when we solve Problem (3.1) using each algorithm. Here, $S$ is assumed to be a compact set of diameter $D$. Bregman distances are utilized in the algorithms of [6] and [55], but we show the number of iterations when the Bregman distance is Euclidean. For the algorithm in [55], the inequality $G(\bar{x})\leq\epsilon$ holds in expectation. In addition, the algorithm in [55] yields $\bar{x}\in S$ such that $f(\bar{x})\leq f^{*}+\frac{3}{7}\epsilon$, while Algorithm 1 yields $\bar{x}\in S$ such that $f(\bar{x})\leq f^{*}+\frac{1}{1+\alpha}\epsilon$. The number of iterations in the algorithm of [6] is obtained by taking the limit as $\beta\to 1$, where $\beta<1$ is a parameter appearing in [6, Corollary 2.1].
@@ 4.1 General copositive programming problems
with the coefficient matrix $\bm{Q}\in\mathcal{S}^{n}$ given by $\bm{A}_{0}+\sum_{i=1}^{m}x_{i}\bm{A}_{i}$. In the following subsubsections, we give an overview of several approaches to solving a standard quadratic programming problem of the form (4.7), both exactly (the case where the violation parameter $\alpha$ equals 0) and inexactly (the case where $\alpha>0$). An exact method is presented in Section 4.1.1. Inexact methods are presented in Sections 4.1.2, 4.1.3, and 4.1.4. Whereas the exact method requires external solvers to solve a mixed-integer linear programming problem, only a finite number of function evaluations are performed in the inexact methods. In Sections 4.1.2, 4.1.3, and 4.1.4, we only consider the case where $\alpha=1$. This is without loss of generality because for general $\alpha>0$, we regard $\alpha\epsilon$ as a new $\epsilon$.
We briefly summarize the three inexact methods introduced in the subsequent subsubsections. The method presented in Section 4.1.2 is deterministic and was originally proposed as a polynomial-time approximation scheme by Bomze and de Klerk [9]. The method in Section 4.1.3 is a brute-force randomized approach. As shown in Section 5.1, this method has limited practical value; however, we include it to motivate the method in Section 4.1.4. The method in Section 4.1.4 is a randomized approach that combines the ideas of Sections 4.1.2 and 4.1.3. When $n$ is large, this method is expected to require fewer function evaluations to approximate $\gamma(\bm{Q})$ than the method in Section 4.1.2.
In the following methods, we frequently use a lower bound for $\gamma(\bm{Q})$. The choice of the lower bound is arbitrary, and there are many alternatives that are summarized in the literature [12]. In our study, because we need to calculate the lower bound per iteration, it might be preferable to adopt a computationally inexpensive one. As such a bound, under the convention that $1/0=\infty$, $a+\infty=\infty$ for any $a\in\mathbb{R}\cup\{\infty\}$, and $1/\infty=0$, Bomze, Locatelli, and Tardella [12] provided the following lower bound for $\gamma(\bm{Q})$:
$\underline{\gamma}(\bm{Q})\coloneqq\min_{1\leq i\leq j\leq n}Q_{ij}+\frac{1}{\displaystyle\sum_{k=1}^{n}\frac{1}{Q_{kk}-{\displaystyle\min_{1\leq i\leq j\leq n}Q_{ij}}}}.$
They showed that this bound outperforms other simple closed-form lower bounds reported in [12, Section 2].
@@ 4.1.1 Exact method through mixed-integer linear programming
To solve the standard quadratic programming problem in (4.7) exactly, we can utilize the mixed-integer linear programming reformulations proposed by Gondzio and Yıldırım [23]. In their paper, various mixed-integer linear programming reformulations of standard quadratic programming problems are presented. Among them, we adopt the following reformulation based on [23, Proposition 2]:*1*1*1Using their terminology, we adopt formulation (MILP2) with the lower bound $\ell_{1}(\bm{Q})$. They claim that this reformulation is suitable for large-scale instances because of its robust practical performance [23, page 319].
@@
Note that finding an optimal solution of the standard quadratic programming problem in (4.7) is more demanding than merely testing copositivity of the matrix $\bm{Q}$. Anstreicher [3] proposed a mixed-integer linear programming formulation for testing copositivity of a symmetric matrix. However, this formulation is not directly connected to standard quadratic programming problems. Therefore, the mixed-integer programming formulation in (4.9) is more suitable for computing an optimal solution of (4.7).
@@ 4.1.2 Inexact deterministic method through a regular grid of the standard simplex
for a positive integer $r$. From [9, Theorem 3.2] (see also [44, Theorem 2]), we have
@@
which is polynomial in $n$ if we regard $r$ as a constant. If the lower bound $\underline{\gamma}(\bm{Q})$ is weak, the number $\lvert\Delta_{r}^{n-1}\rvert$ of evaluations can become prohibitively large. In such cases, rather than using the closed-form bound, it may be preferable to use the bounds obtained by solving semidefinite programming problems as presented in [12].
@@ 4.1.3 Inexact randomized method through uniform sampling from the standard simplex
The second inexact method for solving the standard quadratic programming problem is to discretize the standard simplex $\Delta^{n-1}$ by uniformly sampling from it. Let $\bm{\delta}_{1},\dots,\bm{\delta}_{M}$ be a collection of independent and identically distributed random vectors uniformly distributed on $\Delta^{n-1}$. In Proposition 4.2 shown below, we provide a lower bound for the number $M$ of samples sufficient to ensure that
$\min_{1\leq i\leq M}\bm{\delta}_{i}^{\top}\bm{Q}\bm{\delta}_{i}-\gamma(\bm{Q})\leq\epsilon$ (4.14)
holds with probability at least $1-\phi$ for a given $\phi\in(0,1)$. Throughout this subsubsection, $P$ denotes the uniform distribution on $\Delta^{n-1}$.
@@
By [30, Lemma 6.21], the function $\operatorname*{vol}(B(\bm{x},r)\cap\Delta^{n-1})^{n-1}$ with respect to $\bm{x}\in\Delta^{n-1}$ is concave, so it achieves the minimum at an extreme point of $\Delta^{n-1}$. By the symmetry of $\Delta^{n-1}$, it does so at $\bm{x}=\bm{e}_{1}$.
@@
Below, we present the main result of this subsubsection. For notational convenience, we define
$m(\rho,\phi)\coloneqq\frac{\log\phi}{\log(1-\rho)}$ (4.17)
for $\rho,\phi\in(0,1)$.
@@
Let $\bm{\delta}_{1},\dots,\bm{\delta}_{M}$ be a collection of independent and identically distributed random vectors uniformly distributed on $\Delta^{n-1}$. In addition, let $m$ denote the function defined in (4.17). For $\epsilon\in(0,\sqrt{2}K(\bm{Q})]$ and $\phi\in(0,1)$, if the number $M$ of samples satisfies
@@ 4.2 Testing complete positivity of a matrix
where we use [28, Equation (6.16)] to derive the first inequality and we use (4.33) to derive the second inequality. Using this inequality, the Cauchy–Schwarz inequality, and $\lVert\bm{C}\rVert_{\rm F}=1$, we have
@@
iterations are necessary to test non-complete positivity by using Theorem 4.8. Note, however, that $t$ may need to be chosen large in order to detect non-complete positivity, in which case the number of iterations in (4.34) also becomes large. We revisit the choice of $t$ in Section 5.3.
@@ 4.3 The case of symmetric cones beyond nonnegative orthants
In what follows, we explore a potential approach to deriving a global solution to Problem (4.35). It follows from [19, Section 3] that the optimal value of (4.35) is equal to that of the following copositive programming problem with a scalar variable:
@@ 5.1 Comparison among the methods for solving standard quadratic programming problems
Table 2: The results of solving standard quadratic programming problems by using MILP, Grid, and SUnif. The average objective values obtained by solving standard quadratic programming problems with each method are listed in the “Obj. val.” column and the average computational time required to solve them is listed in the “Time [s]” column. The “Dev. from MILP” column denotes the difference between each average objective value and that obtained by using MILP for the same $n$. The “MM” column denotes the average number $M$ of points $\bm{\delta}_{1},\dots,\bm{\delta}_{M}\in\Delta^{n-1}$ that are used to calculate an approximate optimal value $\min_{1\leq i\leq M}\bm{\delta}_{i}^{\top}\bm{Q}\bm{\delta}_{i}$ for a standard quadratic programming problem of the form (4.7). The symbol >>realmax in the “MM” column indicates that the average number of points exceeds realmax in MATLAB, which is the largest finite floating-point number in IEEE double precision (approximately $1.7977\text{\times}{10}^{308}$).
@@
0.1 $9.9\text{\times}{10}^{16}$ >>3600
SUnif 1 $2.8\text{\times}{10}^{55}$ >>3600
0.1 $2.8\text{\times}{10}^{104}$ >>3600
@@
0.1 $2.5\text{\times}{10}^{22}$ >>3600
SUnif 1 $1.0\text{\times}{10}^{126}$ >>3600
0.1 $1.0\text{\times}{10}^{225}$ >>3600
@@
0.1 $5.7\text{\times}{10}^{35}$ >>3600
SUnif 1 >>realmax >>3600
0.1 >>realmax >>3600
@@
0.1 $5.0\text{\times}{10}^{41}$ >>3600
SUnif 1 >>realmax >>3600
0.1 >>realmax >>3600
@@
Second, to see the effectiveness of the randomized method introduced in Section 4.1.4 (denoted GUnif), we compare it with the other randomized method, SUnif, in terms of solution accuracy and computational time. Recall that in SUnif we sample from the standard simplex $\Delta^{n-1}$, whereas in GUnif we sample points from the regular grid $\Delta_{r}^{n-1}$ for some positive integer $r$. For each $n$, we solved only one of the ten instances. We set $\epsilon$ to $1$. For each $(n,\epsilon)$ we set $r$ as in the previous experiment and choose the number $M$ of samples to be the smallest integer greater than or equal to $c\lvert\Delta_{r}^{n-1}\rvert$, where $c\in\{$5\text{\times}{10}^{-1}$,$1\text{\times}{10}^{-1}$,$1\text{\times}{10}^{-2}$\$\}. We sampled $M$ points $\bm{\delta}_{1},\dots,\bm{\delta}_{M}$ from the standard simplex $\Delta^{n-1}$ when we used SUnif, whereas we sampled from the grid $\Delta_{r}^{n-1}$ when we used GUnif. Then we calculated the value $\min_{1\leq i\leq M}\bm{\delta}_{i}^{\top}\bm{Q}\bm{\delta}_{i}$. Strictly speaking, our choice of $M$ does not follow the results in Sections 4.1.3 and 4.1.4. Here, we compare sampling from the standard simplex with sampling from the regular grid by matching the number of samples.
@@
Tables 3 shows the objective values obtained by solving standard quadratic programming problems with each method and the computational time required to solve them. Here, we present only the results for the case $\epsilon=1$; similar results were also obtained for $\epsilon=0.1$. We did not display the results of $n=5,10$ because in these cases the numbers of points were always less than $50$.
We observed that the objective values obtained by using GUnif were always better than those obtained with SUnif. However, in many cases, the computational time for GUnif was longer than that for SUnif. The MATLAB Profiler showed that our implementation of GUnif spent most time in the built-in function sort, which we used to implement uniform sampling from the regular grid. A more efficient implementation of GUnif would shorten its computational time.
5.2 Solving general copositive programming problems